Determine the value of r so the line that passes through (4,-7) and (-2,r) has a slope of 8/3
Answer by Nate(3500)

the equation slope(m) is: m=((first y value - second y value)/(first x value - second x value)
for the ordered pair (4, -7) and (-2, r) and slope of 8/3
m=((first y value - second y value)/(first x value - second x value)
8/3=((-7-r)/(4-(-2))
8/3=(-7-r)/(6)
16=-7-r
23=-r
r=-23